• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 00:01:23 on October 7,2009

    Entry number 293559

    keyword=Swing shift summary

    Shift crew:
    Carlos Munoz Camacho
    Zafar Ahmed
    Chiranjib Dutta

    16:00 Shift started in production mode.

    18:30 Target alarms, heater was at zero, cooler flow increased.

    19:30 Request MCC 10uA (by closing slits) for Compton studies.

    20:00 By Compton-ers' request, we asked for 50 uA.
    After that, a few more low energy current runs at different values for Compton studies.

    21:20 Back to 100 uA.

    22.05 DAQ problem: ROC28 not responding and unable to reboot. Running without ROC28 for the moment.

    22.37 ROC28 successfully rebooted (next run we will be back to normal Happex configuration).

    00:00 Shift ended running production.